OpenStudy (anonymous):

michele Given the function f(x) = 0.3(4)^x, what is the value of f–1^(6)?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (daniel14):

f-1

OpenStudy (daniel14):

i think

OpenStudy (anonymous):

is answer this michele? 2.312

OpenStudy (michele_laino):

here we have: \[\Large {f^{ - 1}}\left( x \right) = \frac{{{{\log }_{10}}\left( {\frac{{10x}}{3}} \right)}}{{{{\log }_{10}}4}}\] so we get: \[\Large {f^{ - 1}}\left( 6 \right) = \frac{{{{\log }_{10}}\left( {\frac{{10 \times 6}}{3}} \right)}}{{{{\log }_{10}}4}} = \frac{{{{\log }_{10}}\left( {20} \right)}}{{{{\log }_{10}}4}} = 2.16\]

OpenStudy (anonymous):

good thanks michele!

OpenStudy (michele_laino):

thanks! @sallyfield888